Use the following information to calculate the additional funds needed for working capital:

Revenues are $90,000 in the current year

Next year's revenues are expected to equal $100,000

Receivables are usually 20%

Inventory is usually 10% of revenues

Payables are usually 5% of revenues

a. $25,000 b. $22,500 c. $3,000 d. $2,500
Answer #1

Correct answer is (d) $2,500

18,000−20,000 = +2,000
9,000-10,000 = +1,000
4,500-5,000 = -500